OpenStudy (anonymous):

What is the value of x? 4x - 9 = 27

OpenStudy (zkoqmosii):

4x - 9 = 27 + 9 4x = 36 36 / 4 x = 9

OpenStudy (anonymous):

thanks

OpenStudy (zkoqmosii):

Yup ;)

OpenStudy (anonymous):

What is the value of x? 6( - 3 - x) - 2x = 14

OpenStudy (zkoqmosii):

Solve for x: 6 (-3-x)-2 x = 14 Distribute 6 over -3-x. 6 (-3-x) = -18-6 x: -18-6 x-2 x = 14 Combine like terms in -2 x-6 x-18. -6 x-2 x = -8 x: -8 x-18 = 14 Isolate terms with x to the left hand side. Add 18 to both sides: (18-18)-8 x = 14+18 Look for two terms that sum to zero. 18-18 = 0: -8 x = 14+18 Evaluate 14+18. 14+18 = 32: -8 x = 32 Divide both sides by a constant to simplify the equation. Divide both sides of -8 x = 32 by -8: (-8 x)/(-8) = 32/(-8) Any nonzero number divided by itself is one. (-8)/(-8) = 1: x = 32/(-8) Reduce 32/(-8) to lowest terms. Start by finding the GCD of 32 and -8. The gcd of 32 and -8 is 8, so 32/(-8) = (8×4)/(8 (-1)) = 8/8×4/(-1) = 4/(-1): x = 4/(-1) Simplify the sign of 4/(-1). Multiply numerator and denominator of 4/(-1) by -1: Answer: | | x = -4

OpenStudy (anonymous):

thanks